OpenStudy (bieber896):

what is the simplified form of each expression? 7x^-8*6x^3 a. 42/x^5 b. 1/42x^5 c. 42x^11 d. 13x^-5

OpenStudy (bieber896):

@Anaise

OpenStudy (bieber896):

@whpalmer4

OpenStudy (whpalmer4):

\[7x^{-8}*6x^3 = \] remember that we multiply the numbers/coefficients, and we add the exponents of the matching variables: \[3b^2*4c^5 = (3*4)b^{2+5} = 12b^7\]

OpenStudy (bieber896):

ok

OpenStudy (whpalmer4):

Now the other thing you need to remember is that \[x^{-n} = \frac{1}{x^n}\]

OpenStudy (bieber896):

alright so am i supposed to plug in some numbers?

OpenStudy (whpalmer4):

\[7x^{-5}*6x^3\]what is \(7*6\)? what is \(x^{-5}*x^3\)?

OpenStudy (bieber896):

is it 42/x^5?

OpenStudy (bieber896):

@whpalmer4

OpenStudy (whpalmer4):

it is indeed!

OpenStudy (bieber896):

yay! do you have time for 2 more?

OpenStudy (whpalmer4):

\[7x^{-8}*6x^3 = 7*6*x^{-8}*x^3 = 42*x^{-8+3} = 42x^{-5} = \frac{42}{x^5}\]

OpenStudy (whpalmer4):

Post them as new questions, I have to go but others will help if I cannot...

OpenStudy (bieber896):

aww alright